helen has $1.35 in her bank in nickels and dimes. there are 9 more nickels than dimes. find the number she has of each coin.
(can you help me with this problem?)

Answer by checkley75(3666)

N=D+9
10D+5N=1.35
NOW REPLACE N WITH (D+9) & SOLVE FOR D.
.10D+.05(D+9)=1.35
.10D+.05D+.45=1.35 [[THE .45 IS THE PRODUCT OF .05*9=.45]]
.15D=1.35-.45
.15D=.90
D=-.90/.15
D=6 NUMBER OF DIMES
N=6+9
N=15 NUMBER OF NUCKLES.
PROOF
.10*6+.05*15=1.35
.60+.75=1.35
1.35=1.35